OpenStudy (anonymous):

Find the perimeter of a square if one side measures 11 centimeters. Include correct units with your solution.

OpenStudy (anonymous):

You just posted one of these, the solution method is identical

OpenStudy (anonymous):

l*w*^2?

OpenStudy (anonymous):

but this the perimeter not the area

OpenStudy (anonymous):

U don't know what the perimeter is?

OpenStudy (anonymous):

no..:(

OpenStudy (anonymous):

wouldnt that be l^2 *w^2? = the perimeter?

OpenStudy (anonymous):

the perimeter is the total length all around a plane object.

OpenStudy (anonymous):

So for a triangle it would be sum of the lengths of the sides.

OpenStudy (anonymous):

So what is the perimeter here?

OpenStudy (anonymous):

this is for a square. so 11^2*11^2=14641

OpenStudy (anonymous):

44 cm

OpenStudy (anonymous):

oh so it would be 11^4 ?

OpenStudy (anonymous):

I am amazed and astounded sritama, you can calculate a perimeter.....

OpenStudy (anonymous):

i too!!!

OpenStudy (anonymous):

You have a side of 11 and there are 4 of them all the same in a square. As i said, you want the sum of these.

OpenStudy (anonymous):

oh ok... so it would be 11*4 =44

OpenStudy (anonymous):

yes.u r in class?

OpenStudy (anonymous):

yes i am in college math. for the 2nd time .. :( failed by 1 point

OpenStudy (anonymous):

ok